the length of a rectangle is 1 inch less than 3 times the width. if the length is increased by 6 inches and the width is increased by 5, then the length will be twice the width. what are the dimensions of the rectangle
Answer by Finavon(81)

let l & b be length and breadth
1+l=3*b so l=3b-1
l+6=2*(b+5)
substitute for l
3b-1+6=2*(b+5)=2b+10
subtract (2b-5) from each side
b=5
l=3*5-1=14
rectangle is 5 in by 14 in
Check: 3*5 - 1 =14
14+6 = 20
(5+5)*2 = 20